OpenStudy (anonymous):

A wheel of diameter 50.0 cm starts from rest and rotates with a constant angular acceleration of 2.00 rad/s^2. At the instant the wheel has computed its second revolution, compute the radial acceleration of a point on the rim in two ways. (1)Using the relationship . arad=w2r (2)from the relationship arad=v2/r

OpenStudy (anonymous):

arad=w2r so Radial Acceleration = (Angular Velocity)^2 * radius and arad=v2/r Radial Acceleration = (tangential velocity)^2 / radius can you figure out what the Angular Velocity would be if theta = 4pi and using the rotational kinematic equations? with alpha = 2rad/s^2 ? w^2 = w^2 + alpha*theta [kinematic equation i'd suggest to use]